Shift report 20011004 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Thu Oct 04 2001 - 15:49:53 EDT

  • Next message: Apache: "Shift report 20011005 16:00-24:00"

    ------------------------------------------------------------
    Date: Thu Oct  4 15:49:52 2001, Shift: 08:00-16:00
    Supervisor: Eun-Joo Kim, with: 
    Log book#9 pages 89 - 90
    
    Shift Summary:
    ------------------------------------------------------------
    1. Beam for 8:00 - 12:00, about 4 hrs
       "Physics operation for RHIC will resume shortly"
    
    
    2. Run 5126, 5127 :  D1-D5 OFF (MRS@60, FS@3)
                         211k, 65k trigger 1
       Run 5128 :  A-polarity, 83k trigger 1
                   ZDC rate ~ 50-60
       Run 5129, 5130, 5131 : Pedestal, DVM Run
    
    
    3. No new detector problems
       Mainframe FFS1440 : still not connected
                           H1,C1 - checking by "supermon"
       H2 slat #3 : couldn't turn on HV.
       Magnets had problems to turn on, and were fixed
       by magnet person. Now OK.          
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Thu Oct 04 2001 - 15:50:58 EDT